# Matchings in multipartite hypergraphs

### v1

@@ Abstract.
CB gratefully acknowledges support from ERC Advanced Grant 883810, ERC Starting Grant 947978 and Philip Leverhulme Prize PLP–2020–183; RM gratefully acknowledges support from EPSRC Standard Grant EP/R034389/1.
@@ 1.1. Wider context
Kühn and Osthus [25] gave asymptotically best-possible sufficient minimum codegree conditions for a perfect matching in both settings, specifically $\delta^{*}_{k-1}(H)>n/2+O(\sqrt{n\log(n)})$ in the multipartite setting (with $n$ being the common size of the vertex classes) and $\delta_{k-1}(H)>n/2+O(\sqrt{n\log(n)})$ in the non-partite setting (where $n=|V(H)|$ and we assume the necessary condition that $k$ divides $n$ without further comment). For the general setting Rödl, Ruciński and Szemerédi [35] improved this result to the exact best-possible minimum codegree condition for large values of $n$, namely $\delta_{k-1}(H)\geq n/2-k+c$ where $c\in\{1/2,1,3/2,2\}$ depends on the values of $n$ and $k$. Similarly, in the multipartite setting Aharoni, Georgakopoulos and Sprüssel [1] showed that (for large $n$) if $H$ is a $k$-partite $k$-graph such that $\delta_{[k]\setminus\{1\}}(H)>n/2$ and $\delta_{[k]\setminus\{2\}}(H)\geq n/2$, then (with no further conditions on the minimum multipartite codegree into parts $V_{3},\ldots,V_{k}$) $H$ must contain a perfect matching; in particular this implies that the condition $\delta^{*}_{k-1}(H)>n/2$ is a best-possible sufficient condition for the existence of a perfect matching in $H$.
@@ 1.2. Main ideas
Outline proof of Theorem 1.7. To illustrate how the result for rainbow matchings described above can be used to prove Theorem 1.7, we consider here the the case of Theorem 1.7 in which $\sum_{i\in[k]}a_{i}=n$; the argument for the other cases not covered by Fact 1.5 is broadly similar. In this case we wish to find a perfect matching in $H$. Write $q:=\sum_{2\leq i\leq k}a_{i}=n-a_{1}$, so the assumption of Theorem 1.7 implies that $q$ is small. We choose an arbitrary perfect matching $M$ in the complete $k$-partite $k$-graph on $V_{2},\dots,V_{k}$, and use a double-counting argument (Proposition 3.2) to show that there is a subset $X\subseteq V_{1}$ of size close to $n-q$ such that every vertex in $X$ forms an edge with many $(k-1)$-tuples in $M$. Let $Z=V_{1}\setminus X$ be the set of the remaining vertices of $V_{1}$, so $|Z|$ is slightly larger than $q$.
@@ 2. Matchings under a combination of degree and multiplicity
The fact that $H_{j}$ has no dominating set of size at most $(1+2k\varepsilon)q$ then implies that $V\setminus U_{j}$ is not a dominating set in $H_{j}$, and therefore we may choose an edge $e_{j}\in H_{j}[U_{j}]$ as claimed.
@@
$|X|<\left(|D|-|A|\right)\cdot\frac{n^{k-1}}{2}+|A|\cdot\beta n^{k-1}=\frac{|D|n^{k-1}}{2}+(\beta-1/2)|A|n^{k-1}\leq\frac{(1+\mu)qn^{k-1}}{2}+(\beta-1/2)|A|n^{k-1}.$
@@
Observe that for each $x\in X$ there are at most $3\mu q$ vertices $y\in N_{X}$ for which $xy$ is not an edge. So the number of pairs $xy$ with $x\in X$ and $y\in N_{X}$ for which $xy$ is not an edge is at most $3\mu q|X|\leq 4\mu q^{2}$. It follows that at most $20\mu q$ vertices of $N_{X}$ lie in more than $q/5$ such pairs. So we may choose a set $B^{\prime}\subseteq N_{X}\subseteq B$ of size $|B^{\prime}|=\lfloor q/2\rfloor$ not including any such vertex, meaning that every $y\in B^{\prime}$ has at least $|X|-q/5$ neighbours in $X$.
@@
If (1) holds, then arbitrarily choose a perfect rainbow matching $M$ in the family $\{H_{j}:j\in[t]\setminus A\}$. This is possible since $|[t]\setminus A|\leq 2\eta q\leq q/k\leq\max_{i\in[k]}a_{i}$, so we may form $M$ by choosing $t-|A|$ pairwise-disjoint crossing $(k-1)$-tuples which avoid $V_{i^{\prime}}$, where $i^{\prime}:=\max_{i\in[k]}a_{i}$, and greedily extending each to an edge of the respective $H_{j}$ so that no vertex is used more than once. Let $Y:=V(M)$, so $|Y|=k|M|\leq 2k\eta q\leq\mu q$. It follows by (1) that $G[A\cup(V\setminus Y)]$ admits a matching $M_{G}$ which covers the vertices of $A$, which we will shortly use to greedily extend $M$ to a perfect rainbow matching in $\mathcal{H}$.
@@
If instead (2) holds, then set $V_{i}^{\prime}:=V_{i}\setminus Z$ for each $i\in[k]$, $V^{\prime}:=V\setminus Z$, $n^{\prime}:=n-s$, $q^{\prime}=q-s$ and $H^{\prime}_{j}:=H_{j}[V^{\prime}]$ for each $j\in[t]$. Then $H^{\prime}_{1},\dots,H^{\prime}_{t}$ are $k$-partite $k$-graphs with common vertex classes $V^{\prime}_{1},\dots,V^{\prime}_{k}$ each of size at least $n^{\prime}$ such that each crossing $k$-tuple appears in at least $m$ of the graphs $H^{\prime}_{1},\dots,H^{\prime}_{t}$. Moreover, for each $i\in[k]$ and $j\in[t]$ we have $\delta_{[k]\setminus\{i\}}(H_{j})\geq a^{\prime}_{i}:=a_{i}-|Z\cap V_{i}|$, and $\sum_{i\in[k]}a^{\prime}_{i}=\sum_{i\in[k]}a_{i}-|Z|=q^{\prime}$. Since $t\geq m+q=m+q^{\prime}+s\geq m+q^{\prime}+k-1$, we may apply Lemma 2.1 with $[t]\setminus X$ in place of $C$ to obtain a rainbow matching $M$ of size $m+q^{\prime}=m+q-s$ in the family $\{H^{\prime}_{1},\dots,H^{\prime}_{t}\}$ such that $M$ includes an edge of colour $j$ for every $j\in[t]\setminus X$ (i.e. $M$ includes distinct edges for each such $j$). Let $U\subseteq[t]$ be the set of all $j\in[t]$ for which $M$ has no edge of colour $j$, so $U\subseteq X$ and furthermore $|U|=t-(m+q-s)\geq s$. So we may arbitrarily choose a subset $A^{\prime}\subseteq U\subseteq X$ of size $s$, whereupon (2) implies that there is a perfect matching $M_{G}$ in $G[A^{\prime}\cup Z]$, which we will use to extend $M$ to a rainbow matching in $G$ of size $m+q$.
@@
In both cases we now use the matching $M_{G}$ to greedily extend $M$ to a rainbow matching of size $|M|+|M_{G}|$ in $\mathcal{H}$. Observe that each edge $e\in M_{G}$ has the form $j_{e}x_{e}$ for some $j_{e}\in A$ and $x_{e}\in V$. For each edge $e\in M_{G}$ in turn we choose an edge $f_{e}$ of $H_{j_{e}}$ with $x_{e}\in f_{e}$ such that $f_{e}$ does not intersect $V(M)$ or any previously-chosen edge $f_{e^{\prime}}$ and also $f_{e}$ does not contain any vertex $x_{e^{\prime}}$ for $e^{\prime}\neq e$. Observe for this that since $e$ is an edge of $G$ there are at least $\deg_{H_{j_{e}}}(x_{e})\geq n^{k-1}/2$ edges of $H_{j_{e}}$ containing $x_{e}$, and the number of these edges which we are forbidden from using is at most $(|V(M)|+k|M_{G}|)n^{k-2}\leq ktn^{k-2}$, so it is always possible to choose an edge $f_{e}$ with the required properties. The chosen edges $f_{e}$ then form a perfect rainbow matching $M^{\prime}$ of size $|M_{G}|$ in the family $\{H_{j_{e}}:e\in M_{G}\}$, and so $M^{*}:=M\cup M^{\prime}$ is a rainbow matching in $\mathcal{H}$ of size $|M|+|M_{G}|$. It follows that $M^{*}$ is perfect if (1) holds, and has size $m+q$ if (2) holds; either way this completes the proof.
@@
Since $a_{1}\geq n-q-k+3\geq n-(1+\varepsilon)q+1$ we may additionally insist that $|X|\leq a_{1}$ (by removing vertices if necessary). Let $Y:=V_{1}\setminus X$, so $n-a_{1}\leq|Y|\leq(1+\varepsilon)q$ by (i). Applying Proposition 3.1 to $Y$ we find that each crossing $(k-1)$-tuple avoiding $V_{1}$ appears in at least $m:=|Y|-(n-a_{1})\geq 0$ of the link graphs $\{L_{u}(H)\}_{u\in Y}$. Since $q\leq n-a_{1}$ we have $|Y|\geq m+q$, so by Lemma 2.5, applied with $m$ and $q$ as here and with $t=|Y|$, the family $\{L_{u}(H)\}_{u\in Y}$ admits a rainbow matching $M_{rainbow}$ of size $m+q\leq|Y|$. This means there is a set $Z\subseteq Y$ of size $m+q$ and a bijection $\psi:Z\to M_{rainbow}$ such that for each $z\in Z$ we have $\{z\}\cup\psi(z)\in E(H)$.
Let $M^{\prime}\subseteq M$ be the matching in $K^{k-1}$ consisting of all edges of $M$ which do not intersect any edge of $M_{rainbow}$, and let $M^{*}$ be a matching in $K^{k-1}$ of size $n$ with $M^{\prime}\cup M_{rainbow}\subseteq M^{*}$. Since $|M\setminus M^{\prime}|\leq|V(M_{rainbow})|\leq(k-1)|Y|\leq(k-1)(1+\varepsilon)q$, by (ii) we have for each $u\in X$ that
$|M^{*}\cap L_{u}(H)|$ $\geq|M^{\prime}\cap L_{u}(H)|\geq|M\cap L_{u}(H)|-|M\setminus M^{\prime}|\geq(1+\varepsilon)kq-(k-1)(1+\varepsilon)q\geq(1+\varepsilon)q.$
@@ 4. Concluding Remarks
As described in Section 1.2, we see the combination of degree and multiplicity conditions appearing in Lemma 2.5 as being a natural setting to consider, and expect that that result should hold for a much wider range of values of $m,q$ and $t$; we think this question is worthy of further attention.
5. Acknowledgements
@@
For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.

### v2

@@ Abstract.
CB gratefully acknowledges support from ERC Advanced Grant 883810, ERC Starting Grant 947978 and Leverhulme Trust Early Career Fellowship ECF–2023–393; RM gratefully acknowledges support from EPSRC Standard Grant EP/R034389/1.
@@ 1.1. Wider context
[…]
@@ 1.2. Main ideas
Outline proof of Theorem 1.7. To illustrate how the result for rainbow matchings described above can be used to prove Theorem 1.7, we consider here the case of Theorem 1.7 in which $\sum_{i\in[k]}a_{i}=n$; the argument for the other cases not covered by Fact 1.5 is broadly similar. In this case we wish to find a perfect matching in $H$. Write $q:=\sum_{2\leq i\leq k}a_{i}=n-a_{1}$, so the assumption of Theorem 1.7 implies that $q$ is small. We choose an arbitrary perfect matching $M$ in the complete $(k-1)$-partite $(k-1)$-graph on $V_{2},\dots,V_{k}$, and use a double-counting argument (Proposition 3.2) to show that there is a subset $X\subseteq V_{1}$ of size close to $n-q$ such that every vertex in $X$ forms an edge with many $(k-1)$-tuples in $M$. Let $Z=V_{1}\setminus X$ be the set of the remaining vertices of $V_{1}$, so $|Z|$ is slightly larger than $q$.
@@ 2. Matchings under a combination of degree and multiplicity
The fact that $H_{j}$ has no dominating set of size at most $(1+2k\varepsilon)q$ then implies that $V\setminus U_{j}$ is not a dominating set in $H_{j}$, and therefore we may choose an edge $e_{j}\in H_{j}[U_{j}]$ as required for (c).
@@
$|X|$ $<\left(|D|-|A|\right)\cdot\frac{n^{k-1}}{2}+|A|\cdot\beta n^{k-1}=\frac{|D|n^{k-1}}{2}+(\beta-1/2)|A|n^{k-1}$
$\leq\frac{(1+\mu)qn^{k-1}}{2}+(\beta-1/2)|A|n^{k-1}.$
@@
For each $x\in X$, since $|N_{X}|\leq(1+2\mu)q$ and $\deg(x)\geq(1-\mu)q$, Observe that for each $x\in X$ there are at most $3\mu q$ vertices $y\in N_{X}$ for which $xy$ is not an edge. So the number of pairs $xy$ with $x\in X$ and $y\in N_{X}$ for which $xy$ is not an edge is at most $3\mu q|X|\leq 4\mu q^{2}$. It follows that at most $20\mu q$ vertices of $N_{X}$ lie in more than $q/5$ such pairs. So we may choose a set $B^{\prime}\subseteq N_{X}\subseteq B$ of size $|B^{\prime}|=\lfloor q/2\rfloor$ not including any such vertex, meaning that every $y\in B^{\prime}$ has at least $|X|-q/5$ neighbours in $X$.
@@
If (a) holds, then arbitrarily choose a perfect rainbow matching $M$ in the family $\{H_{j}:j\in[t]\setminus A\}$. This is possible since $|[t]\setminus A|\leq 2\eta q\leq q/k\leq\max_{i\in[k]}a_{i}$, so we may form $M$ by choosing $t-|A|$ pairwise-disjoint crossing $(k-1)$-tuples which avoid $V_{i^{\prime}}$, where $i^{\prime}$ is such that $a_{i^{\prime}}:=\max_{i\in[k]}a_{i}$, and greedily extending each to an edge of the respective $H_{j}$ so that no vertex is used more than once. Let $Y:=V(M)$, so $|Y|=k|M|\leq 2k\eta q\leq\mu q$. It follows by (a) that $G[A\cup(V\setminus Y)]$ admits a matching $M_{G}$ which covers the vertices of $A$, which we will shortly use to greedily extend $M$ to a perfect rainbow matching in $\mathcal{H}$.
@@
If instead (b) holds, then set $V_{i}^{\prime}:=V_{i}\setminus Z$ for each $i\in[k]$, $V^{\prime}:=V\setminus Z$, $n^{\prime}:=n-s$, $q^{\prime}=q-s$ and $H^{\prime}_{j}:=H_{j}[V^{\prime}]$ for each $j\in[t]$. Then $H^{\prime}_{1},\dots,H^{\prime}_{t}$ are $k$-partite $k$-graphs with common vertex classes $V^{\prime}_{1},\dots,V^{\prime}_{k}$ each of size at least $n^{\prime}$ such that each crossing $k$-tuple appears in at least $m$ of the graphs $H^{\prime}_{1},\dots,H^{\prime}_{t}$. Moreover, for each $i\in[k]$ and $j\in[t]$ we have $\delta_{[k]\setminus\{i\}}(H_{j})\geq a^{\prime}_{i}:=a_{i}-|Z\cap V_{i}|$, and $\sum_{i\in[k]}a^{\prime}_{i}=\sum_{i\in[k]}a_{i}-|Z|=q^{\prime}$. Since $t\geq m+q=m+q^{\prime}+s\geq m+q^{\prime}+k-1$, we may apply Lemma 2.1 with $[t]\setminus X$ in place of $C$ to obtain a rainbow matching $M$ of size $m+q^{\prime}=m+q-s$ in the family $\{H^{\prime}_{1},\dots,H^{\prime}_{t}\}$ such that $M$ includes an edge of colour $j$ for every $j\in[t]\setminus X$ (i.e. $M$ includes distinct edges for each such $j$). Let $U\subseteq[t]$ be the set of all $j\in[t]$ for which $M$ has no edge of colour $j$, so $U\subseteq X$ and furthermore $|U|=t-(m+q-s)\geq s$. So we may arbitrarily choose a subset $A^{\prime}\subseteq U\subseteq X$ of size $s$, whereupon (b) implies that there is a perfect matching $M_{G}$ in $G[A^{\prime}\cup Z]$, which we will use to extend $M$ to a rainbow matching in $G$ of size $m+q$.
@@
In both cases we now use the matching $M_{G}$ to greedily extend $M$ to a rainbow matching of size $|M|+|M_{G}|$ in $\mathcal{H}$. Observe that each edge $e\in M_{G}$ has the form $j_{e}x_{e}$ for some $j_{e}\in A$ and $x_{e}\in V$. For each edge $e\in M_{G}$ in turn we choose an edge $f_{e}$ of $H_{j_{e}}$ with $x_{e}\in f_{e}$ such that $f_{e}$ does not intersect $V(M)$ or any previously-chosen edge $f_{e^{\prime}}$ and also $f_{e}$ does not contain any vertex $x_{e^{\prime}}$ for $e^{\prime}\neq e$. Observe for this that, since $e$ is an edge of $G$, there are at least $\deg_{H_{j_{e}}}(x_{e})\geq n^{k-1}/2$ edges of $H_{j_{e}}$ containing $x_{e}$, and the number of these edges which we are forbidden from using is at most $(|V(M)|+k|M_{G}|)n^{k-2}\leq ktn^{k-2}$, so it is always possible to choose an edge $f_{e}$ with the required properties. The chosen edges $f_{e}$ then form a perfect rainbow matching $M^{\prime}$ of size $|M_{G}|$ in the family $\{H_{j_{e}}:e\in M_{G}\}$, and so $M^{*}:=M\cup M^{\prime}$ is a rainbow matching in $\mathcal{H}$ of size $|M|+|M_{G}|$. It follows that $M^{*}$ is perfect if (a) holds, and has size $m+q$ if (b) holds; either way this completes the proof.
@@
Since $a_{1}\geq n-q-k+3\geq n-(1+\varepsilon)q+1$ we may additionally insist that $|X|\leq a_{1}$ (by removing vertices from $X$ if necessary). Let $Y:=V_{1}\setminus X$, so $n-a_{1}\leq|Y|\leq(1+\varepsilon)q$ by (a). Applying Proposition 3.1 to $Y$ we find that each crossing $(k-1)$-tuple avoiding $V_{1}$ appears in at least $m:=|Y|-(n-a_{1})\geq 0$ of the link graphs $\{L_{u}(H)\}_{u\in Y}$. Since $q\leq n-a_{1}$ we have $|Y|\geq m+q$, so by Lemma 2.5, applied with $m$ and $q$ as here and with $t=|Y|$, the family $\{L_{u}(H)\}_{u\in Y}$ admits a rainbow matching $M_{rainbow}$ of size $m+q\leq|Y|$. This means there is a set $Z\subseteq Y$ of size $m+q$ and a bijection $\psi:Z\to M_{rainbow}$ such that for each $z\in Z$ we have $\{z\}\cup\psi(z)\in E(H)$.
Let $M^{\prime}\subseteq M$ be the matching in $K^{k-1}$ consisting of all edges of $M$ which do not intersect any edge of $M_{rainbow}$, and let $M^{*}$ be a matching in $K^{k-1}$ of size $n$ with $M^{\prime}\cup M_{rainbow}\subseteq M^{*}$. Since $|M\setminus M^{\prime}|\leq|V(M_{rainbow})|\leq(k-1)|Y|\leq(k-1)(1+\varepsilon)q$, by (b) we have for each $u\in X$ that
$|M^{*}\cap L_{u}(H)|$ $\geq|M^{\prime}\cap L_{u}(H)|\geq|M\cap L_{u}(H)|-|M\setminus M^{\prime}|$
$\geq(1+\varepsilon)kq-(k-1)(1+\varepsilon)q\geq(1+\varepsilon)q.$
@@ 4. Concluding Remarks
As described in Section 1.2, we see the combination of degree and multiplicity conditions appearing in Lemma 2.5 as being a natural setting to consider, and expect that result should hold for a much wider range of values of $m,q$ and $t$; we think this question is worthy of further attention.
Acknowledgements
@@
We thank the anonymous reviewers for their helpful suggestions for improvements to the manuscript.